OpenStudy (anonymous):

brody's farm borders a river. He buys 1,000 meters of fencing. What dimensions maximize the area of the farm?

OpenStudy (anonymous):

First write what you need to find A = LW Then what you know (since it is against a river you only need three sides hence the one W) P = 2L + W 1000 = 2L + W Solve for one variable (I chose W) W = 1000 - 2L Plug into Area equation A = 1000L - 2L^2 Take the derivative dA/dL = 1000 - 4L Solve for L L=250 Plug into perimeter equation 1000 = 2(250) + W W = 500 Therefore the dimensions should be 250 m x 500 m given that it is against the river.

OpenStudy (anonymous):

you got this answer from yahoo

OpenStudy (anonymous):

no I figure it out by myself did I get it right or not